A modern penny weighs 2.3850 g but contains only 0.0600 g of copper (Cu). How many copper atoms are present in a modern penny? x 10 Cu atoms Report your answer using scientifie notation.
Answer #1

Answer

mass of cooper in modern penny is 0.0600 g

finding the moles of cooper in 0.600 g by the formula moles=mass/molar mass

moles of cooper = 0.0600/63.546 = 9.44*10^-4 moles of cooper

number of atoms= moles of cooper *avagadro number

number of atoms of Cu= 9.44*10^-4*6.02*10^23 = 5.68*10^20 atoms of Cu